OpenStudy (lillian_a):

c(d) = 3√(5d^2 − 1)^5

OpenStudy (lillian_a):

\[\sqrt[3]{(5d^{2}-1)^{5}}\] Find the derivative in the simplest factored form

OpenStudy (solomonzelman):

Oh derivatives ... nice

OpenStudy (solomonzelman):

Ok, what woud be the derivative of \(\large\color{#000000 }{ \displaystyle \sqrt[3]{x} }\) ?

OpenStudy (solomonzelman):

HINT: \(\large\color{#000000 }{ \displaystyle \sqrt[3]{x} =x^{1/3} }\)

OpenStudy (lillian_a):

\[x^{-2/3}\]

OpenStudy (solomonzelman):

you forgot something

OpenStudy (solomonzelman):

The power rule: \(\large\color{#000000 }{ \displaystyle \frac{d}{dx} x^n=\color{red}{n} x^{n-1}}\)

OpenStudy (solomonzelman):

you woud have to multiply times 1/3

OpenStudy (lillian_a):

oh yeah

OpenStudy (solomonzelman):

The power rule: \(\large\color{#000000 }{ \displaystyle \frac{d}{dx} x^{1/3}=\color{red}{(1/3) } x^{-2/3}}\)

OpenStudy (solomonzelman):

Woud you mind if we use, just: \(\large\color{#000000 }{ \displaystyle f(x)=\sqrt[3]{(5x^{2}-1)^{5}} }\) because to write derivatives as d/dd is a bit wierd ??

OpenStudy (solomonzelman):

lets use x instead if you don;t mind

OpenStudy (lillian_a):

okay

OpenStudy (lillian_a):

i think i figured it out

OpenStudy (solomonzelman):

now do the chain rue

OpenStudy (solomonzelman):

chain rule**

OpenStudy (solomonzelman):

\(\large\color{#000000 }{ \displaystyle f(x)=\sqrt[3]{(5x^{2}-1)^{5}} }\) You know that, \(\large\color{#000000 }{ \displaystyle \frac{d}{dx} \sqrt[3]{x}=\frac{1}{3}x^{-2/3}}\) But inside the cube root you have another function of x. (Right?) So you need to apply the chain rule, as follows: \(\large\color{#000000 }{ \displaystyle f(x)=\sqrt[3]{(5x^{2}-1)^{5}} }\) \(\large\color{#000000 }{ \displaystyle f'(x)=\frac{1}{3} \left[(5x^{2}-1)^{5}\right]^{-2/3} \color{blue}{\times \frac{d}{dx} (5x^{2}-1)^{5}} }\)

OpenStudy (solomonzelman):

have you seen the chain rue previously?

OpenStudy (lillian_a):

yeah

OpenStudy (solomonzelman):

So, do you understand everything we have done so far?

OpenStudy (lillian_a):

yup , I think i figured it out

OpenStudy (solomonzelman):

Ok, now you need to differentiate \(\large\color{#000000 }{ \displaystyle \color{blue}{ (5x^{2}-1)^{5}} }\)

OpenStudy (solomonzelman):

((You will need anothe Chain Rule))

OpenStudy (solomonzelman):

What would be the derivative of x^5 ?

OpenStudy (lillian_a):

5x^4

OpenStudy (solomonzelman):

Yes, but since instead of: \(\large\color{#000000 }{ \displaystyle \color{blue}{ (\color{red}{x})^{5}} }\) You have, \(\large\color{#000000 }{ \displaystyle \color{blue}{ (\color{red}{5x^{2}-1})^{5}} }\) You therefore must apply a chain rule (the second time)

OpenStudy (solomonzelman):

What is the derivative of 5x^2-1 ?

OpenStudy (lillian_a):

10x

OpenStudy (solomonzelman):

yes very good.

OpenStudy (solomonzelman):

Now I will post the entire work for the derivative.. I wil try to be as clear and brief as I can

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle f(x)=\sqrt[3]{(5x^{2}-1)^{5}} }\) we know that, \(\color{#000000 }{ \displaystyle \frac{d}{dx} \sqrt[3]{x}=\frac{1}{3}x^{-2/3}}\) But inside the cube root you have another function of x. So we apply the Chain Rule (in blue), as follows: \(\color{#000000 }{ \displaystyle f'(x)=\frac{1}{3} \left[(5x^{2}-1)^{5}\right]^{-2/3} \color{blue}{\times \frac{d}{dx} (5x^{2}-1)^{5}} }\) we know that, \(\color{#000000 }{ \displaystyle \frac{d}{dx} x^5=5x^4 }\) But we have a function of x raised to the 5th power (not just x), and therefore the expression inside the 5th power needs a Chain Rule (in green). \(\color{#000000 }{ \displaystyle f'(x)=\frac{1}{3} \left[(5x^{2}-1)^{5}\right]^{-2/3} \color{blue}{\times 5(5x^{2}-1)^{4}} \color{green}{\times \frac{d}{dx} \left(5x^2 -1\right)}}\) We know that: \(\color{#000000 }{ \displaystyle \frac{d}{dx} (5x^2-1)=2\cdot5x+0= 10x }\) So we can write the derivative as: \(\color{#000000 }{ \displaystyle f'(x)=\frac{1}{3} \left[(5x^{2}-1)^{5}\right]^{-2/3} \color{blue}{\times 5(5x^{2}-1)^{4}} \color{green}{\times 10x}}\)

OpenStudy (solomonzelman):

Now all is needed is to simplify using algebra.

OpenStudy (solomonzelman):

I wil post properties that might be helpful ...

OpenStudy (solomonzelman):

\(\color{#bb4400}{ {\rm [1]}\quad\quad \displaystyle {\bf y}^{-A}= \frac{1}{{\bf y}^{A}} \quad\quad\Longrightarrow \quad\quad \displaystyle {\bf y}^{-N/M}= \frac{1}{{\bf y}^{N/M}} }\) \(\color{#bb4400}{ {\rm [2]}\quad\quad \displaystyle {\bf y}^{Q} \times{\bf y}^{R} ={\bf y}^{Q+R} }\)